A gas has a volume of 28 liters at STP. Calculate the pressure if the temperature is increased to 35°C and the volume is decreased to 15 liters


1
Expert's answer
2022-05-13T01:00:02-0400

V1=28 l=0.028 m3

T1=273 K

P1=101325 Pa

V2=15 l=0.015 m3

T2=35 °C=308 K

P2-?

R=8.31 J/(mole K)


PV=nRT

"n=\frac{P_1V_1}{RT_1}=\frac{101325x0.028}{8.31x273}=1.25 mole"

"P_2=\frac{nRT_2}{V_2}=\frac{1.25x8.31x308}{0.015}=213389 Pa"